find the general solution of the differential equation using undetermined coefficients, y^(iv) -2y"'+2y"=3e^-x+2e^-x.x+e^-sinx.


1
Expert's answer
2021-03-01T07:20:44-0500

find the general solution of the differential equation using undetermined coefficients, y^(iv) -2y"'+2y"=3e^-x+2e^-x.x+e^-sinx.

Solution:

We can solve differential equation using undetermined coefficients if the right (heterogeneous) part of the equation is in the form:

"e^nP_m" , where "P_m" is a polynomial of degree "m";

or in the form:

"e^\alpha(P_k\cos{(\beta x)}+P_l\sin{(\beta x)})" , where "P_k" and "P_l" are polynomials of degrees "k" and "l".

Since "e^{-\sin{x}}" cannot be presented as "e^nP_m" or "e^\alpha(P_k\cos{(\beta x)}+P_l\sin{(\beta x)})", we can conclude that a question consists mistake. Assume the correct equation is the following:

"y^{IV}-2y'''+2y''=3e^{-x}+2e^{-x}\cdot x+e^{-x}\cdot \sin{x}" (*)

Let's find the complementary solution to this differential equation:

"y^{IV}-2y'''+2y''=0"

The characteristic equation for this differential equation is the following:

"r^4-2r^3+2r^2=0"

"r^2(r^2-2r+2)=0"

"r=0" ; "r=1\pm i"

The complementary solution:

"y_c=C_1+C_2\cdot x+e^{x}(C_3\cos{x}+C_4\sin{x})"

Let's proceed with finding a particular solution. The form of the particular solution would be:

"y_p=e^{-x}(Ax+B+C\cos{x}+D\sin{x})"

"y_p'=e^{-x}""(A-Ax-B+(D-C)\cos{x}-(C+D)\sin{x})"

"y_p''=e^{-x}""(Ax-2A+B-2D\cos{x}+2C\sin{x})"

"y_p'''=e^{-x}""(-Ax+3A-B+2(C+D)\cos{x}+2(D-C)\sin{x})"

"y_p^{IV}=e^{-x}""(Ax-4A+B-4C\cos{x}-4D\sin{x})"

Substituting "y_p^{IV}", "y_p'''" and "y_p''" into the (*) gives:

"e^{-x}(Ax-4A+B-4C\cos{x}-4D\sin{x})-""2e^{-x}(-Ax+3A-B+2(C+D)\cos{x}" "+2(D-C)\sin{x})+" "2e^{-x}(Ax-2A+B-2D\cos{x}+2C\sin{x})=" "3e^{-x}+2e^{-x}\cdot x+e^{-x}\cdot \sin{x}"

Choose "A" , "B" , "C", and "D" so that the coefficients of the exponentials on either side of the equal sign are the same:

"A=\frac25" , "B=\frac{43}{25}" , "C=\frac{1}{16}", "D=-\frac{1}{16}".

The particular solution:

"y_p=e^{-x}(\frac25x+\frac{43}{25}+\frac{1}{16}\cos{x}-\frac{1}{16}\sin{x})"

General solution:

"y=y_c+y_p=" "C_1+C_2\cdot x+e^{x}(C_3\cos{x}+C_4\sin{x})+" "e^{-x}(\frac25x+\frac{43}{25}+\frac{1}{16}\cos{x}-\frac{1}{16}\sin{x})"

Solution:

"y=C_1+C_2\cdot x+e^{x}(C_3\cos{x}+C_4\sin{x})+" "e^{-x}(\frac25x+\frac{43}{25}+\frac{1}{16}\cos{x}-\frac{1}{16}\sin{x})"